(a) CFor the following reaction at equilibrium at 2000°C, the concentration of N 2 and O are both
5.2 M.

N 2 (g) + O 2 (g) ↔ 2 NO 2 (g) K eq = 6.2 × 10 -4

Calculate the concentration of NO at equilibrium.
(b) Acetic acid, HC 2 H 3 O 2 , is in equilibrium with its ions:

HC 2 H 3 O 2 (aq) ↔ H + (aq) + C 2 H 3 O 2 – (aq) K eq = 1.8 × 10 -5

At equilibrium, the concentration of the ions are:
[H + ] = 1.33 × 10 -3 M
[C 2 H 3 O 2 – ] = 1.33 × 10 -3 M
Calculate the concentration of the acid, HC 2 H 3 O
1
Expert's answer
2020-06-15T14:43:25-0400

a) [NO] = ([O2] × [N2] × K)1/2 = (5.2 × 5.2 × 6.2 × 10-4)1/2 = 0.129 M

b) [C2H4O2] = [H+] × [C2H3O2-] / K = 1.33 × 10-3 × 1.33 × 10-3 / (1.8 × 10-5) = 0.098 M